OpenStudy (anonymous):

I NEED HELP PLEASE I WILL MEDAL AND FAN!!! Identify which table shows a direct variation. A. both tables B. Table 1 only C. Table 2 only D. neither table The graph is in the comments below.

OpenStudy (anonymous):

Is there a picture?

OpenStudy (anonymous):

You mean B

OpenStudy (anonymous):

I dont like direct answers you need to help me to understand the problem

OpenStudy (anonymous):

I'll tag people cuz i'm bad at graphs. @texaschic101 @welshfella

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Nnesha @perl @Spyro517

OpenStudy (anonymous):

hey i gtg bye

OpenStudy (welshfella):

a direct variation ihas the equation y = kx where k is some number so k = y/x and if the values of y/x are constant then you know you have dirsct variation

OpenStudy (welshfella):

Table 1 work out 7.5 / 3 11.25/4.5 17.5 / 7 and see if they give the same result

OpenStudy (welshfella):

if they d then table 1 shows direct variation

OpenStudy (welshfella):

* they do

OpenStudy (welshfella):

perform similar calculations for Table 2 3 / 1.5 = 12/8 = 19.5/13 =

OpenStudy (anonymous):

So yes? @welshfella or @texaschic101

OpenStudy (welshfella):

Hi did you do those calculations?

OpenStudy (anonymous):

I don't know because to be honest I don't know how to divide with decimals because I am terrible at math.

OpenStudy (welshfella):

use ur calculator 7.5, press divide button, 3 , then press '=' or 'exe'

OpenStudy (anonymous):

Yes they all end up with a direct variation.

OpenStudy (anonymous):

Right @welshfella?

OpenStudy (welshfella):

table 1 you mean?

OpenStudy (anonymous):

Yes sorry.

OpenStudy (welshfella):

yes table 1 is direct variation because the results all come to 2.5

OpenStudy (welshfella):

Try Table 2 calculations

OpenStudy (anonymous):

So that means it is B.?

OpenStudy (anonymous):

Correct?

OpenStudy (welshfella):

if only table 1 is direct variation it is B but have you tested Table 2?

OpenStudy (anonymous):

Yes and all of the results came to 1.5...so that means it is A?

OpenStudy (welshfella):

try Table 2 again

OpenStudy (anonymous):

Ok

OpenStudy (welshfella):

3 / 1.5 = 12 / 8 = 19.5 / 13 =

OpenStudy (anonymous):

@welshfella I got 3/1.5=2, 12/8=1.5, and 19.5/13=1.5

OpenStudy (welshfella):

right - so is this direct variation?

OpenStudy (welshfella):

are the 3 results all the same?

OpenStudy (anonymous):

No so my answer is B.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

?

OpenStudy (welshfella):

right Table 1 only B

OpenStudy (anonymous):

Thank you.

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

I got more I'll post another one sec.....